OpenStudy (anonymous):

find the domain of the composite function f o g f(x) = x+4; g(x)= 9/x+6

OpenStudy (anonymous):

i think it is x not equal to -6

OpenStudy (amistre64):

the domain of the function f(g(x)) doesnt change any from it component parts. f(x) can use anything ; but g(x) cant use -6 so the domain of the composite is everything but "-6"

OpenStudy (anonymous):

yay my answer is right:)

OpenStudy (amistre64):

g(f(x)) would give us an additional ixnay for x... in that we might not be able to use -10

OpenStudy (anonymous):

yes ..in the case of g o f

OpenStudy (anonymous):

it is x not equal to 10

OpenStudy (amistre64):

if we were to consider functions that would perhaps"cancel"out an inapproriate domain, then the original domains would still take precident

OpenStudy (amistre64):

your gonna tell me im wrong..aintcha :)

OpenStudy (anonymous):

so ?? what say tamas?

OpenStudy (anonymous):

If \[f(x)=x+4\textrm{ and }g(x)=\frac{9}{x+6}\] then \[(f\circ g)(x)=f(g(x))=\frac{9}{x+6}+4.\] So \[D_{f\circ g}=\mathbb{R}\backslash\{-6\}.\]

OpenStudy (anonymous):

I concur with gorbe.tamas.